# Pritam Goswami    Adri Bhattacharya    Raja Das    Partha Sarathi Mandal

### v1

@@ 5.2.1 Correctness and Complexity
Let $a_{i},a_{j}$ and $a_{k}$ be three agents that starts from two initial nodes, say $home_{1}$ and $home_{2}$. By Pigeon hole principle, exactly one of $home_{1}$ and $home_{2}$ initially must have two agents. Without loss of generality let $home_{1}$ has two agents, say $a_{i}$ and $a_{k}$, initially. In this case, the agents having multiplicity greater than one at the current node does not move. On the other hand the singleton agent, i.e., $a_{j}$ starting from $home_{2}$ moves clockwise marking each node with message right. If $a_{j}$ reaches $home_{1}$ before being consumed by the black hole, $home_{1}$ now has three agents co-located. Thus from here the agents execute the whiteboard version of PerpExplore-Coloc-Pbl (refer Remark 4.24). On the other hand, let us consider the case when $a_{j}$ gets consumed before reaching $home_{1}$. Note that irrespective of the location of $home_{2}$ it takes at most $n-1$ rounds for $a_{j}$ to reach $home_{1}$ from the beginning. So $a_{i}$ and $a_{j}$ waits for $n$ rounds and finds that no one has arrived yet. In this case, both of $a_{i}$ and $a_{k}$ moves to $home_{2}$ along clockwise direction and starts to perform the cautious walk , where lowest ID agent among $a_{i}$ and $a_{k}$ changes its state to Cautious-Leader, whereas the other agent changes its state to Cautious-Follower. Next, the agent executing Cautious-Leader searches for the marking right. As argued earlier, within at most $4n$ rounds after an agent is consumed, at least one of the remaining alive agents can detect the black hole and continue to explore the ring avoiding that node.
So from Remark 4.24 and from the results in this section we have the following theorem
Theorem 5.24.
Three synchronous agents are necessary and sufficient to solve the problem PerpExploration-BBH on a ring $R$, when each node of $R$ has a whiteboard of $O(\log n)$ bits of memory irrespective of their starting location.
6 Conclusion

### v2

@@ 11.2 Correctness and Complexity
Let $a_{i},a_{j}$, and $a_{k}$ be three agents that start from two initial nodes, say $home_{1}$ and $home_{2}$. By Pigeon hole principle, exactly one of $home_{1}$ and $home_{2}$ initially must have two agents. Without loss of generality let $home_{1}$ have two agents, say $a_{i}$ and $a_{k}$, initially. In this case, the agents having multiplicity greater than one at the current node do not move. On the other hand the singleton agent, i.e., $a_{j}$ starting from $home_{2}$ moves clockwise marking each node with message right. If $a_{j}$ reaches $home_{1}$ before being destroyed by the Byzantine black hole, $home_{1}$ now has three agents co-located. Thus from here the agents execute the whiteboard version of PerpExplore-Coloc-Pbl (refer Subsection 4.3). On the other hand, let us consider the case when $a_{j}$ gets destroyed before reaching $home_{1}$. Note that irrespective of the location of $home_{2}$ it takes at most $n-1$ rounds for $a_{j}$ to reach $home_{1}$ from the beginning. So $a_{i}$ and $a_{j}$ waits for $n$ rounds and finds that no one has arrived yet. In this case, both $a_{i}$ and $a_{k}$ move to $home_{2}$ along the clockwise direction and start to perform the cautious walk, where the lowest ID agent among $a_{i}$ and $a_{k}$ changes its state to Cautious-Leader, whereas the other agent changes its state to Cautious-Follower. Next, the agent executing Cautious-Leader searches for the marking right. As argued earlier, within at most $4n$ rounds after an agent is destroyed, at least one of the remaining alive agents, detect the Byzantine black hole and continue to explore the ring avoiding that node.